# ReMatch: Retrieval Enhanced Schema Matching with LLMs

### v1

@@ Abstract.
Schema matching is a crucial task in data integration, involving the alignment of a source database schema with a target schema to establish correspondence between their elements. This task is challenging due to textual and semantic heterogeneity, as well as differences in schema sizes. Although machine-learning-based solutions have been explored in numerous studies, they often suffer from low accuracy, require manual mapping of the schemas for model training, or need access to source schema data which might be unavailable due to privacy concerns. In this paper we present a novel method, named ReMatch, for matching schemas using retrieval-enhanced Large Language Models (LLMs). Our method avoids the need for predefined mapping, any model training, or access to data in the source database. In the ReMatch method the tables of the target schema and the attributes of the source schema are first represented as structured passage-based documents. For each source attribute document, we retrieve $J$ documents, representing target schema tables, according to their semantic relevance. Subsequently, we create a prompt for every source table, comprising all its attributes and their descriptions, alongside all attributes from the set of top $J$ target tables retrieved previously. We employ LLMs using this prompt for the matching task, yielding a ranked list of $K$ potential matches for each source attribute. Our experimental results on large real-world schemas demonstrate that ReMatch significantly improves matching capabilities and outperforms other machine learning approaches. By eliminating the requirement for training data, ReMatch becomes a viable solution for real-world scenarios.
@@ 1. Introduction
Schema matching is a fundamental task in data management and integration, involving the identification of semantic correspondences between elements of two or more database schemas (Madhavan et al., 2001). This process is essential as it lays the groundwork for various data manipulation and integration tasks, including data warehousing, database federation, and the merging of information systems. At its core, schema matching aims to establish mappings between schema elements that are semantically related, regardless of differences in naming, structure, or data type.
The necessity for schema matching arises from the ever-growing volume of data generated by diverse applications and organizations, where data is often trapped in siloed repositories, each with its unique schema. In domains ranging from healthcare to retail, efficient schema matching can lead to more informed decision-making, seamless integration of heterogeneous systems, and ultimately, a competitive advantage in data-driven insights.
However, the schema matching task is challenging due to several inherent complications. Firstly, schemas are designed with different perspectives and terminologies, namely textual heterogeneity, reflecting the conceptualization of domain experts from disparate fields. This semantic heterogeneity can lead to ambiguous mappings where schema elements have the same name but different meanings, or different names but the same meaning. Secondly, structural heterogeneity compounds this complexity, with schemas exhibiting varied architectures, hierarchies, constraints, and model granularity differences.
@@
Human schema matching, a manual and time-consuming process, requires significant effort from skilled individuals. This can be expensive and impractical, particularly in large-scale projects. Furthermore, human matchers are prone to errors and inconsistencies due to cognitive biases and fatigue (Shraga et al., 2021). Their performance can also be influenced by the complexity and ambiguity of the schema elements, leading to potential inaccuracies in the matched results. As a result, the automation of schema matching has become a major focus within the AI and database-oriented research community over the years.
@@
In recent years large language models (LLMs) have achieved significant advancements across many challenging tasks that require a deep understanding of semantics. This includes tasks that until recently required significant human efforts. These models have shown an impressive ability to generalize to new tasks without any task-specific fine-tuning, even in areas significantly divergent from the ones they were originally trained on, including various data related tasks (Narayan et al., 2022; Zhang et al., 2023a).
@@
In this paper, we present a new approach that unlike previously proposed machine learning (ML) methods, circumvents the need for predefined mapping, model training, or access to data in the source database. Our approach, which we call ReMatch, improves the task of schema matching by using retrieval-enhanced LLMs. Fig. 1 illustrates our proposed approach. First, we represent schema tables and attributes (columns) as structured documents (1); then, for each source attribute, we retrieve semantically similar target tables (2); Lastly, we create a prompt containing source attributes and target candidate attributes, and employ LLM for the task of selecting top $K$ most relevant target attributes, yielding a list of potential $K$ matches for each attribute in the source schema (3).
@@
Our contributions in this study are as follows: (1) Introducing a new method for the task of schema matching, which allows for scalable and accurate matching results, without any model training or access to labeled data. (2) Proposing a mechanism to reduce the search space of the target schema for efficient candidate generation, inspired by techniques used in information retrieval (IR), and in particular, by retrieval augmented generation (RAG) methods. (3) Exploiting the generative abilities of LLMs, and their text comprehension to perform semantic ranking between two schemas, instead of Cartesian-product-based classification, in alignment with human matchers. (4) Demonstrating the effectiveness of our proposed method, in the challenging healthcare domain, and provide a complete mapping between two important healthcare schemas. To the best of our knowledge, this is currently the largest publicly available schema matching dataset.
The rest of the paper is structured as follows. The next section provides background on LLMs and on schema matching. Section 3 presents our proposed method. Section 4 describes the experimental evaluation and results. Section 5 reviews related work, Section 6 discusses insights from the obtained results and limitations, and finally, Section 7 concludes the paper.
2. Background
2.1. Large Language Models (LLMs)
A paradigm shift in language models (LMs) was initiated with the introduction of Transformer models by Vaswani et al.(Vaswani et al., 2017). This architecture enables models to capture long-range dependencies between input and output sequences, overcoming limitations in previous models. The power of transformer models is reflected in generative and embedding models such as the GPT model family (Radford et al., 2018, 2019; Brown et al., 2020) and BERT (Devlin et al., 2019). Following the introduction of ChatGPT (OpenAI, 2022), the terms Large Language Model and LLMs have become synonymous with general purpose pretrained language models.
@@ 2.1. Large Language Models (LLMs)
Generative LMs like GPT are trained to generate human-like text and can be fine-tuned for a variety of tasks, but also can be reused for a variety of tasks, with no additional training (Radford et al., 2018; Brown et al., 2020; Liu et al., 2023; Mirchandani et al., 2023). Embedding models like BERT or Ada (Neelakantan et al., 2022) provide contextual embeddings that have significantly improved performance on a wide range of NLP tasks (Sun et al., 2019; Minaee et al., 2021). Moreover, the representation of text via embeddings allows for efficient and accurate passage retrieval, using semantic similarity (Neelakantan et al., 2022; Lewis et al., 2020).
2.2. Schema Matching
Schema matching is a crucial process in various domains such as data warehousing, E-commerce, and semantic web services. It involves identifying correspondences between different data schemas and creating a mapping between them. A schema, in its simplest definition, is a representation of the structure of a database. Schema matching, therefore, plays a vital role in data integration, enabling the consolidation of various data sources into a unified, coherent form.
@@ 2.2. Schema Matching
Traditionally, schema matching was executed manually, requiring significant human effort. However, this approach strongly relies on the expertise of human matchers (Dragisic et al., 2016). This has led to development of automated schema matching, primarily using ML algorithms (Do and Rahm, 2002; Bernstein et al., 2011; Madhavan et al., 2001). Although automatic, previous work on ML-based schema matching was limited to very basic schemas, with most works dealing only with toy datasets (Do and Rahm, 2002; Gal et al., 2021; Mudgal et al., 2018). This stemmed from a combination of lack of high-quality datasets (Koutras et al., 2020), as well as relatively poor results of past methods.
Recently, following the advancements in NLP, mechanisms with enhanced understanding of language semantics have been able to achieve better results. Leveraging LLMs for schema matching has been shown to provide initial promising results, principally due to their ability to understand and interpret the semantics and context of the data schemas (Li et al., 2020; Zhang et al., 2021).
@@ 3. ReMatch
In this section, we present our new approach, called ReMatch, inspired by techniques from IR in general, and by LLMs in particular, for the task of schema matching. The basic steps of our approach are illustrated in Fig. 1. Below, we provide further details regarding the concepts underlying our overall methodology, as well as pseudo code of relevant algorithms.
3.1. Problem Statement
Given two database schemas, a source schema $\mathcal{S}_{1}$ and a target schema $\mathcal{S}_{2}$, with sets of tables $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ and sets of attributes $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively, the schema matching task involves finding a mapping between $(\mathcal{A}_{1},\mathcal{T}_{1})\in\mathcal{S}_{1}$ to $(\mathcal{A}_{2},\mathcal{T}_{2})\in\mathcal{S}_{2}$.
The matches between their attributes are captured by a relation match: $\mathcal{P}(\mathcal{A}_{1})\times\mathcal{P}(\mathcal{A}_{2})$. An element $(A_{1},A_{2})\in$ match defines a matching pair possibly representing the same information in the schemas. If $\lvert A_{1}\rvert=1$ and $\lvert A_{2}\rvert=1$, we call the match an elementary match or 1:1 match. Otherwise, we refer to it as a complex match or m:n match.
The goal of automatic schema matching is to find all those matches that are meaningful. Usually, there are conditions that a collection of meaningful matches has to obey, for instance, that the matches do not overlap. That is, for every pair of matches $(a_{1},a_{2})$ and $(a_{3},a_{4})$ in the mapping it both holds that $a_{1}\cap a_{3}=\emptyset$ and $a_{2}\cap a_{4}=\emptyset$. We call this collection of matches a mapping. From a structural point of view, a mapping is an element of the power set of matches.
Formally, we would like to find a function $\Psi:\mathcal{A}_{1}\rightarrow\mathcal{P}(\mathcal{A}_{2})$, such that $\forall a\in\mathcal{A}_{1},\>\forall a^{\prime}\in\Psi(a),\>(a,a^{\prime})\in match$. In other words, $\Psi(a)=A^{\prime}\subseteq\mathcal{A}_{2}$, should contain all relevant matches for $a$.
@@ 3.1. Problem Statement
$\mathcal{C}_{s}\leftarrow\emptyset$, $\mathcal{C}_{t}\leftarrow\emptyset$, $\Psi_{K}\leftarrow\{\}$ /* lines 2-4 are performed once, as a pre-processing step. */
1 $\mathcal{C}_{t}\leftarrow TableToDoc(\mathcal{S}_{2},\mathcal{T}_{2},\mathcal{A}_{2})$ $\mathcal{C}_{s}\leftarrow AttributeToDoc(\mathcal{S}_{1},\mathcal{T}_{1},\mathcal{A}_{1})$ $\tilde{\mathcal{C}_{t}}\leftarrow\Phi(\mathcal{C}_{t})$ foreach $t_{i}\in\mathcal{T}_{1}$ do
2       $\mathcal{T}_{c}\leftarrow\{\}$ foreach $a_{k}\in\mathcal{A}_{1}[t_{i}]$ do
3             $\mathcal{T}_{c}\leftarrow RankTopjTables(\mathcal{C}_{s}[t_{i};a_{k}],\mathcal{C}_{t},\tilde{\mathcal{C}_{t}},\Phi,J)$
@@ 2       $\mathcal{T}_{c}\leftarrow\{\}$ foreach $a_{k}\in\mathcal{A}_{1}[t_{i}]$ do
Inputs : a source attribute as a document $\mathcal{C}_{s}[t_{i};a_{k}]$; the corpus of target tables as documents and its embedding $\mathcal{C}_{t}$ and $\tilde{\mathcal{C}_{t}}$; an embedding model $\Phi$; $J\in\mathbb{N}$.
Outputs : a list of $J$ retrieved target tables.
$\tilde{a}_{emb}\leftarrow\Phi(\mathcal{C}_{s}[t_{i};a_{k}])$ $scores\leftarrow Similarity(\tilde{a}_{emb},\tilde{\mathcal{C}_{t}})$ $top_{j}\leftarrow\mathrm{argmaxes}(\{scores_{1},\ldots,scores_{n}\})[1:j]$ return: $\mathcal{C}_{t}[top_{j}]$
Algorithm 2 RankTopjTables
3.2. Method Description
The proposed methodology leverages textual descriptions, tables and attributes constraints, and data types, for creating an alignment between source and target schemas. To this end, we look at the problem from an IR perspective. We adapt passage-based modeling such that each schema element is represented as a structured document, containing all its textual descriptions and information. We utilize this representation to judge the similarity and retrieve relevant documents. We then derive correspondences based on LLM generation, by producing prompts containing source schema attributes (columns) with relevant target schema candidates.
The methodology is composed of three stages: Given a source schema $\mathcal{S}_{1}$ with a set of tables $\mathcal{T}_{1}$ and a set of attributes $\mathcal{A}_{1}$, and a target schema $\mathcal{S}_{2}$ with a set of tables $\mathcal{T}_{2}$ and a set of attributes $\mathcal{A}_{2}$:
  1. (1)
    Target schema tables and source schema attributes are first transformed into two corpora of structured documents, $\mathcal{C}_{t}$ and $\mathcal{C}_{s}$, respectively. Each target schema table and source schema attribute is represented as a structured document consisting of four descriptive paragraphs. The title of the document is the table’s name, and the opening paragraph provides an overview of the table’s purpose and characteristics. The subsequent paragraphs detail the set of attributes serving as the table’s primary key, the set of attributes referring to other tables (foreign keys), and the rest of the attributes belonging to this table, respectively. Each attribute is followed by its data type and a textual description. For attribute documents, the specific attribute is highlighted above the title.
  2. (2)
    Given constraints on prompt size, we employ a retrieval strategy to allow for scalability to large schemas. For each attribute in the source schema, we search for the top $J$ documents that represent candidate tables from the target schema. These tables should contain the most promising candidate attributes for deriving correspondences. This step ensures that only the most relevant documents are considered for matching, thereby coping with the limitations imposed by the maximal prompt size. To facilitate the retrieval of candidate tables, we utilize a text embedding model to encode both the candidate source attribute, accompanied by its table description, and the corpus of target table documents. These embeddings serve as a basis for measuring semantic similarity, enabling the efficient retrieval of candidate tables. Finally, for every source table $t_{i}\in\mathcal{T}_{1}$, we create a set of all top $J$ candidate tables $\mathcal{T}_{c}$ retrieved.
  3. (3)
    In the last step, the LLM is tasked with selecting the top $K$ most similar target attributes from the set of retrieved tables $\mathcal{T}_{c}$ identified earlier. The model assesses the similarity based on the context provided by the document representations, yielding a ranked list of $K$ potential matches for each attribute in the source schema.
An example of a full prompt, including the outputs of stage (1) and (2) is given in Appendix AAlgorithm 1 presents the entire ReMatch method, where the functions TableToDoc and AttributeToDoc refer to the process from stage (1), and Algorithm 2 focuses on the document ranking process from stage (2).
@@ 4.1. Dataset Creation
MIMIC-III to OMOP  For the first dataset, we created a mapping between the schema of MIMIC-III (Johnson et al., 2016) and The Observational Medical Outcomes Partnership Common Data Model (OMOP) [^1] https://www.ohdsi.org/data-standardization/. MIMIC-III is a public database containing deidentified records from patients who were admitted to the critical care units of the Beth Israel Deaconess Medical Center. OMOP is an open-source data standard, created to standardize the structure and content of healthcare related data. This scenario closely resembles real-world use cases, where a proprietary database created for a specific purpose (a source schema) is mapped to a given industry standard (a target schema) for further uses.
The mapping was created manually by a domain expert, aided by the mapping created in (Paris et al., 2021). The full MIMIC-III schema was mapped. If no matching attribute in OMOP could be found, the attribute was assigned NA. For convenience we will refer to this dataset as MIMIC.
@@
OMAP Benchmark Synthea Dataset  Following (Narayan et al., 2022) we used the Synthea (Walonoski et al., 2018) to OMOP dataset from the OMAP benchmark (Zhang et al., 2021). The dataset contains a partial mapping of the schema for Synthea (the source), a synthetic healthcare dataset and data generator, to a partial subset of relevant OMOP attributes (the target). For convenience we will refer to this dataset as Synthea.
@@
It is important to note that although both datasets are similar in their goals, the first maps the entire MIMIC-III schema to OMOP, while Synthea contains only partial mappings. To the best of our knowledge, the MIMIC dataset is the largest single-schema to single-schema dataset published. Table 1 contains statistics detailing both datasets. The full MIMIC dataset is publicly available at https://github.com/meniData1/MIMIC_2_OMOP.

### v2

@@ Abstract.
Schema matching is a crucial task in data integration, involving the alignment of a source schema with a target schema to establish correspondence between their elements. This task is challenging due to textual and semantic heterogeneity, as well as differences in schema sizes. Although machine-learning-based solutions have been explored in numerous studies, they often suffer from low accuracy, require manual mapping of the schemas for model training, or need access to source schema data which might be unavailable due to privacy concerns. In this paper we present a novel method, named ReMatch, for matching schemas using retrieval-enhanced Large Language Models (LLMs). Our method avoids the need for predefined mapping, any model training, or access to data in the source database. Our experimental results on large real-world schemas demonstrate that ReMatch is an effective matcher. By eliminating the requirement for training data, ReMatch becomes a viable solution for real-world scenarios.
Schema Matching, Large Language Models (LLMs), RAG, Machine Learning, Ranking, Data Integration
ccs: Information systems Mediators and data integrationccs: Computing methodologies Natural language processing
@@ 1. Introduction
Schema matching is a fundamental task in data management and integration, involving the identification of semantic correspondences between elements of two or more database schemas (Madhavan et al., 2001), regardless of differences in naming, structure, or data type. The matching task is challenging since schemas are designed using different perspectives and terminologies. This semantic heterogeneity can lead to ambiguous mappings where schema elements have the same name but different meanings, or different names but the same meaning.
@@
Human schema matching, a manual and time-consuming process, requires significant effort from skilled individuals. This can be expensive and impractical, particularly in large-scale projects. Furthermore, human matchers are prone to errors and inconsistencies due to cognitive biases and fatigue (Shraga et al., 2021). As a result, the automation of schema matching has become a major focus within the AI and database-oriented research community over the years.
@@
In recent years, LLMs have achieved significant advancements across many challenging tasks that require a deep understanding of semantics. These models have shown an impressive ability to generalize to new tasks without any task-specific fine-tuning, even in areas significantly divergent from the ones they were originally trained on, including various data related tasks (Narayan et al., 2022; Zhang et al., 2023a).
@@
In this paper, we present a new approach that unlike previously proposed machine learning (ML) methods, circumvents the need for predefined mapping, model training, or access to data in the source database. Our approach, which we call ReMatch, improves the task of schema matching by using retrieval-enhanced LLMs. Fig. 1 illustrates our proposed approach.
@@
Our contributions in this study are as follows: (1) Introducing a new method for the task of schema matching, which allows for scalable and accurate matching results, without any model training or access to labeled data. (2) Proposing a mechanism to reduce the search space of the target schema for efficient candidate generation. (3) Exploiting the generative abilities of LLMs to perform semantic ranking between two schemas, in alignment with human matchers. (4) Providing a complete mapping between two important healthcare schemas, aiming to address the lack of real-world evaluation datasets in the field. To the best of our knowledge, this is currently the largest publicly available schema matching dataset.
2. Background and Related Work
@@
Large Language Models (LLMs)  Generative language models are trained to generate human-like text and can be fine-tuned for a variety of tasks, but also can be reused for a variety of tasks, with no additional training (Radford et al., 2018; Brown et al., 2020; Liu et al., 2023; Mirchandani et al., 2023). Embedding models like BERT or Ada (Neelakantan et al., 2022) provide contextual embeddings that have significantly improved performance on a wide range of NLP tasks (Sun et al., 2019; Minaee et al., 2021). Moreover, the representation of text via embeddings allows for efficient and accurate passage retrieval, using semantic similarity (Neelakantan et al., 2022; Lewis et al., 2020).
LLMs have been applied to various data preprocessing tasks, including error detection, data imputation, entity matching, as well as schema matching (Mirchandani et al., 2023; Narayan et al., 2022). These models have shown promising results, however, these approaches still faced challenges related to computational expense and inefficiency, as well as lacking evaluation on data that represents real-world schema matching challenges.
@@
Schema Matching  Schema matching involves identifying correspondences between different data schemas. Traditionally, schema matching was done manually, however, this approach strongly relies on the expertise of human matchers (Dragisic et al., 2016). This has led to development of automated schema matching, primarily using ML algorithms (Do and Rahm, 2002; Bernstein et al., 2011; Madhavan et al., 2001). Much of the work on ML-based schema matching was limited to basic schemas, with most works dealing only with toy datasets (Do and Rahm, 2002; Gal et al., 2021; Mudgal et al., 2018; Shraga et al., 2020; Li et al., 2020), or focused primarily on entity matching only. This stemmed from a combination of lack of high-quality datasets (Koutras et al., 2020), as well as relatively poor results of past methods.
Recent work, such as SMAT (Zhang et al., 2021) and LSM (Zhang et al., 2023b), leverages advancements in natural language processing to achieve semantic mappings between source and target schemas. While these methods improve previous results, they still require extensive data tagging, limiting their practicality in real-world applications.
Finally, the majority of previous methods dealt with schema-matching as a binary classification task over $\mathcal{A}_{1}\times\mathcal{A}_{2}$. While simplifying training and evaluation, in normalized schemas positive labels scale by $O(n)$, whereas negative labels scale by $O(n^{2})$, which is problematic when inference is expensive, as in the case of LLMs.
@@ 3. ReMatch
Problem Statement  Given a source schema $\mathcal{S}_{1}$ and a target schema $\mathcal{S}_{2}$, with sets of tables $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ and sets of attributes $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, respectively, the schema matching task involves finding a mapping between $(\mathcal{A}_{1},\mathcal{T}_{1})\in\mathcal{S}_{1}$ to $(\mathcal{A}_{2},\mathcal{T}_{2})\in\mathcal{S}_{2}$. The matches between their attributes are captured by a relation match: $\mathcal{P}(\mathcal{A}_{1})\times\mathcal{P}(\mathcal{A}_{2})$. An element $(A_{1},A_{2})\in$ match defines a matching pair possibly representing the same information in the schemas. If $\lvert A_{1}\rvert=1$ and $\lvert A_{2}\rvert=1$, we call the match an elementary match or 1:1 match. Otherwise, we refer to it as a complex match or m:n match.
Formally, we want to find a function $\Psi:\mathcal{A}_{1}\rightarrow\mathcal{P}(\mathcal{A}_{2})$, such that $\forall a\in\mathcal{A}_{1},\>\forall a^{\prime}\in\Psi(a),\>(a,a^{\prime})\in match$. In other words, $\Psi(a)=A^{\prime}\subseteq\mathcal{A}_{2}$, should contain all relevant matches for $a$.
@@
Method Description   Algorithm 1 shows the pseudocode for our method. Given a source schema $\mathcal{S}_{1}$ with a set of tables $\mathcal{T}_{1}$ and a set of attributes $\mathcal{A}_{1}$, and a target schema $\mathcal{S}_{2}$ with a set of tables $\mathcal{T}_{2}$ and a set of attributes $\mathcal{A}_{2}$:
(1) Target schema tables and source schema attributes are transformed into two corpora of structured documents, $\mathcal{C}_{t}$ and $\mathcal{C}_{s}$, respectively. Each target schema table and source schema attribute is represented as a structured document consisting of descriptive paragraphs. The title of the document is the table’s name, and the opening paragraph provides an overview of the table’s purpose and characteristics. The subsequent paragraphs detail the set of attributes serving as the table’s primary key, the set of attributes referring to other tables (foreign keys), and the rest of the attributes belonging to this table, respectively. Each attribute is followed by its data type and a textual description. In the case of attribute documents, the specific attribute is highlighted above the title.
(2) For each attribute in the source schema, we search for the top $J$ documents that represent candidate tables from the target schema. This step ensures that only the most relevant documents are considered for matching. To facilitate the retrieval of candidate tables, we utilize a text embedding model to encode each source attribute’s document, and the corpus of target table documents. These embeddings serve as a basis for measuring semantic similarity, enabling the retrieval of candidate tables. Finally, for every source table $t_{i}\in\mathcal{T}_{1}$, we create a set of all top $J$ candidate tables $\mathcal{T}_{c}$ retrieved.
(3) In the last step, the LLM is tasked with selecting the top $K$ most similar target attributes from the set of retrieved tables $\mathcal{T}_{c}$ identified earlier. The model assesses the context provided by the document representations, and yields a ranked list of $K$ potential matches for each attribute in the source schema.
@@
1 $\mathcal{C}_{s}\leftarrow\emptyset$, $\mathcal{C}_{t}\leftarrow\emptyset$, $\Psi_{K}\leftarrow\{\}$
/* lines 2-4 are performed once, as a pre-processing step. */
2 $\mathcal{C}_{t}\leftarrow TableToDoc(\mathcal{S}_{2},\mathcal{T}_{2},\mathcal{A}_{2})$
3 $\mathcal{C}_{s}\leftarrow AttributeToDoc(\mathcal{S}_{1},\mathcal{T}_{1},\mathcal{A}_{1})$
4 $\tilde{\mathcal{C}_{t}}\leftarrow\Phi(\mathcal{C}_{t})$
5 foreach $t_{i}\in\mathcal{T}_{1}$ do
6       $\mathcal{T}_{c}\leftarrow\{\}$
7       foreach $a_{k}\in\mathcal{A}_{1}[t_{i}]$ do
8             $\tilde{a}_{emb}\leftarrow\Phi(\mathcal{C}_{s}[t_{i};a_{k}])$
9             $scores\leftarrow Similarity(\tilde{a}_{emb},\tilde{\mathcal{C}_{t}})$
10             $top_{j}\leftarrow\mathrm{argmaxes}(\{scores_{1},\ldots,scores_{n}\})[1:j]$
11             $\mathcal{T}_{c}\leftarrow\mathcal{T}_{c}\cup\mathcal{C}_{t}[top_{j}]$
@@ 4.1. Dataset Creation
MIMIC-III to OMOP  For the first dataset, we created a mapping between the schema of MIMIC-III (Johnson et al., 2016), a public database containing deidentified records from the Beth Israel Deaconess Medical Center, and The Observational Medical Outcomes Partnership Common Data Model (OMOP) [^1] https://www.ohdsi.org/data-standardization/, an open-source data standard. The mapping was created manually by a domain expert, aided by the mapping created in (Paris et al., 2021). If no matching attribute in OMOP could be found, the attribute was assigned NA. For convenience we will refer to this dataset as MIMIC.
@@
OMAP Benchmark Synthea Dataset  Following (Narayan et al., 2022) we used the Synthea (Walonoski et al., 2018) to OMOP dataset from the OMAP benchmark (Zhang et al., 2021). The dataset contains a partial mapping of the schema for Synthea (the source), a synthetic healthcare dataset, to a partial subset of relevant OMOP attributes (the target). For convenience we will refer to this dataset as Synthea.
@@
It is important to note that although both datasets are similar, the first maps the entire MIMIC-III schema to OMOP, while Synthea contains only partial mappings. To the best of our knowledge, the MIMIC dataset is the largest single-schema to single-schema dataset published. Table 1 contains statistics detailing both datasets. The full MIMIC dataset is publicly available at https://github.com/meniData1/MIMIC_2_OMOP.
4.2. Experiments
We established the performance of ReMatch and SMAT (Zhang et al., 2021), the previous state-of-the-art (SOTA) method, on MIMIC and Synthea.
As our method treats schema matching as a retrieval problem, the most appropriate metric is accuracy@$K$, as defined in Eq. 1. Unlike the F1-score, standard accuracy@$K$ is not well defined for m:n matches. To deal with this, we limited ourselves to evaluating 1:1 or m:1 relations.  [^2] It is worth noting that for a dataset with only 1:1 and m:1 mappings, accuracy@1 is equivalent to F1-score when using argmax, instead of a threshold, for prediction (since it trivially reflects both the precision and the recall simultaneously).
There are multiple dimensions we would like to optimize. The first is maximizing accuracy@$K$. We also seek lower values of $K$, as our aim is to assist human schema matching, and minimize cognitive overload (Bernstein et al., 2011; Chernev et al., 2015). Finally, we would like to minimize the number of target tables, reducing prompt size and costs.
Evaluating ReMatch  For all experiments we used GPT-4 (OpenAI, 2023), specifically GPT-4-1106, and Ada2 (Greene et al., 2022; Neelakantan et al., 2022) as our embedding model.
We performed a grid search over various values of $J$ and $K$, also allowing us to select hyperparameters for further evaluations. As an ablation study, we evaluated the performance on MIMIC when only the names of the tables and attributes are used for retrieval and generation. We also evaluated a version with no retrieval stage, i.e., for each table in the source schema we used the entire target schema. Finally, we ran ReMatch multiple times on each of the setups in the ablation study, to verify stability across generations.
Evaluating ReMatch with Guidance  A human matcher may determine that the suggestions are not good enough, or decide to provide information they already have. In our evaluation we investigated a mechanism, inspired by the relevance feedback technique (Rocchio, 1971). To do so, we selected a mapping of the form $(T_{1},a_{1})\rightarrow(T_{2},a_{2})$ for a single attribute from each table. Specifically, we used the SUBJECT_ID attribute where available (19/25 tables), and an attribute containing unique identifiers otherwise. We then automatically included $T_{2}$ in the set of retrieved tables, $\mathcal{T}_{c}$, and provided the LLM with this mapping in the prompt.
Evaluating SMAT  Our method’s performance was evaluated against SMAT (Zhang et al., 2021), a previously SOTA non-LLM model. For SMAT’s training we used the default hyperparameters: {Learning Rate: 0.8, Batch Size: 64, Epochs: 30}, with early stopping, and 10% of the training data was used as a validation set to avoid over-fitting. Models were trained on a single Nvidia A-100 GPU.
We tried two setups, 80% training and 20% testing, and vice versa. The splits were stratified across table names. We also extended the evaluation to include accuracy@$K$. We calculated the accuracy@$K$ by determining whether the correct match received one of the $K$ highest classification scores. Binary predictions were given by using the optimal threshold for the validation set.